Five students held a party. They agreed to share the expenses equally. Leroy spent $14 on drinks. Alex spent $3 on paper plates. Kulwinder spent $7 on decorations. Max spent $9 on snacks. Bobbi spent $2 on envelopes and paper to send invitations out, and she also spent $5.80 in postage. Who owes money to whom?
Answer by edjones(8007)

14+3+7+9+2+5.8=40.8
40.8/5=8.16
Leroy:14-8.16=5.84
Alex:3-8.16=-5.16
Kulw:7-8.16=-1.16
Max:9-8.16=0.84
Bobbi:7.80-8.16=-.36
.
The negatives owe the positives.
.
Ed